Charges are equal and opposite on two parallel plates. When space was evacuated, the electric field was 2.80 x 105 V/m. When it was filled with a dielectric material, the electric field was 1.20x105 V/m. What is the dielectric constant? What is the material?

1
Expert's answer
2022-04-19T12:06:04-0400
"\kappa=\frac{E}{E_0}=\frac{2.80*10^5}{1.20*10^5}=2.33"